OpenStudy (anonymous):

what do i do with a slope problem like this 3x+2y=6

OpenStudy (anonymous):

You are trying to find the slope?

OpenStudy (amistre64):

Ax +By = C slope = -A/B

OpenStudy (amistre64):

3x +2y = 6 slope = -3/2

OpenStudy (anonymous):

^ yes